OpenStudy (anonymous):

Circle A has an equation of x2 + (y + 3)2 = 16. Where is the center of the circle located? on the x-axis in the third quadrant on the y-axis in the first quadrant

OpenStudy (anonymous):

well i worked on it & got an answer but i'm not sure if it is right would u like to k it ?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok x^2+ (y+3)^2=16 X^2+Y^2+6y+9=16 therefore: -X^2=Y^2+6Y-7 so we need another equ. to solve the 2 variables that what i couldn't reach sorry

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

but thank u

OpenStudy (anonymous):

its on the y axis, since the center is equal to (0,-3)